At what points does the line 4x + 12 = 3y cross the x and y axes?
Found 2 solutions by scott8148, matthewatcfs:
Answer by scott8148(6628)

when the line crosses the x axis, y=0; and vice versa

4(0)+12=3y ... 12=3y ... 4=y ... so y-intercept is (0,4)

4x+12=3(0) ...4x+12=0 ... 4x=-12 ... x=-3 ... so x-intercept is (-3,0)

Answer by matthewatcfs(3)

No problem at all!
Think about it for a moment.
When something crosses the x axis, it has NO y value--or zero. So, anything that intersects the x axis would have the point (#,0).
When something crosses the y axis, it has NO x value--or zero. So, anything that intersects the y axis would have the point (0,#).
Now, using that information, to find where 4x + 12 = 3y crosses the x axis, simply plug 0 in for y--since y has to be 0 for it to be true.
4x + 12 = 3y
4x + 12 = 3(0)
4x + 12 = 0
4x + 12 - 12 = 0 - 12
4x = -12
4x/4 = -12/4
x = -3
This line crosses the x axis at the point (-3,0).
Now, the y axis.
To find where 4x + 12 = 3y crosses the y axis, simply plug 0 in for x--since x has to be 0 for it to be true.
4x + 12 = 3y
4(0) + 12 = 3y
0 + 12 = 3y
12 = 3y
12/3 = 3y/3
4 = y
This line crosses the x axis at the point (0, 4).